# Tame deformations of highly singular function germs

### v2

@@ Abstract.
We give analytic and algebraic conditions under which a deformation of real analytic functions with non-isolated singular locus is locally topologically trivial at the boundary.
@@
deformations, fibrations, equisingularity
@@
The authors acknowledge the partial support of the Labex CEMPI (ANR-11-LABX-0007-01).
@@ 1. Introduction
For the same question in the complex setting, i.e. for holomorphic function germs with isolated singularities, the constancy of the Jacobian ideal $(\partial F_{t})$ obviously implies that the Milnor number $\mu(F_{t})=\dim_{{\mathbb{C}}}\frac{{\mathcal{O}}_{n}}{(\partial F_{t})}$ is constant, and thus the homology of the Milnor fibre is constant too since it is concentrated in the group $\tilde{H}_{n-1}(F_{t},{\mathbb{Z}})$, which is free of rank precisely $\mu(F_{t})$. Moreover, by the Lê-Ramanujam [LR], Timourian [Tim], and King [Ki] results, the constancy of the Milnor number implies that the deformation $F_{t}$ is topologically trivial with the exception of the surface case (i.e. $n=3$) which remains open. In case of nonisolated singularity, this question seems to be open in general. There is nevertheless a rich literature on equisingularity problems for families of functions, dealing with various algebraic-geometric sufficient conditions, starting with the Whitney equisingularity studied by Teissier [Te1, Te2], extended by Gaffney [Ga], Houston [Hou] and many others.
In the real setting there are results by King [Ki] on the topological triviality of one-parameter families with isolated singularity. While the non-isolated case remains widely open in general, Parusinski’s paper [Pa2] goes beyond “isolated singularity” in the setting of families of the form $F(x,t)=f(x)+tg(x)$ subject to the condition:
@@
We will approach Question 1 from a different angle, by focussing on the constancy of the tube fibration in a fixed ball, as explained in the following.
Let $F_{t}(x)=F(x,t)$ be a ${\mathbb{K}}$-analytic deformation of the singular function germ $F_{0}=f:({\mathbb{K}}^{n},0)\to({\mathbb{K}},0)$ having any singular locus, where ${\mathbb{K}}={\mathbb{R}}$ or ${\mathbb{C}}$, and where $F$ is of class C1 in the variable $t$.
@@
We say that the deformation $F$ of the singular function germ $f$ is topologically trivial at the boundary if for any small enough $0<\delta\ll r$ the restriction:
@@
is a locally trivial fibration, which is independent, modulo isotopy, of $r$ and $\delta$, and of the parameter $t$ with $|t|\leq\eta\ll\delta$. (Compare to the local tube fibration, Definition 3.1.)
@@
In the real setting, $D_{\delta}$ denotes an interval centred at the origin and its boundary $\partial D_{\delta}=\{a_{-},a_{+}\}$ consists of two points. In this setting, the independency asked by Definition 1.1 means the invariance of the two fibres [^1] Of which one may be empty, but not both empty since we assume that the function $F_{0}$ is non-constant., $B\cap F_{t}^{-1}(a_{-})$ and $B\cap F_{t}^{-1}(a_{+})$. One may then ask:
@@
Question 2. What natural (and minimal) condition implies that the deformation $F$ is topologically trivial at the boundary?
@@
The problem posed by Question 2 is not a classical equisingularity problem and makes really sense in case of nonisolated singularities. If the function germ $F_{0}$ has an isolated singularity, then for $t\not=0$ this may either split into several isolated singularities of $F_{t}$, or not split at all. Even if the isolated singularity of $F_{0}$ splits, the tube fibrations (2) still exists precisely because the splitting is concentrated at the origin only, see e.g. [JiT] for the study of a more general setting. In case it does not split, then one has stronger results: the above cited ones by Lê-Ramanujam [LR] and Timourian [Tim] over ${\mathbb{C}}$, and by King [Ki] over ${\mathbb{R}}$, yield the topological triviality of the family of function germs $F_{t}$, or what one calls topological equisingularity.
@@
In the setting of non-isolated holomorphic functions, the concept of “topologically trivial at the boundary” (Definition 1.1) appeared under the name of “admissible” deformations in the studies of the topology of certain classes of function germs with non-isolated singular locus, starting with the seminal paper by Siersma [Si1] on “line singularities”, and the follow-up more general studies of the Milnor fibre of function germs with 1-dimensional and 2-dimension singular loci, see e.g. [Si2], [Pe], [dJ], [Za], [MS], [Ne], [Fe], [FM], [ST2] etc. The very recent paper [Hof] gives algebro-geometric criteria which are sufficient to establish admissibility for complex-analytic deformations.
@@
In both settings, complex and real, our general answer to Question 2 is (Theorem 3.7): “tame deformations” are topologically trivial at the boundary. This “tameness” condition, introduced in Definition 3.5, is based on the transversality of all small enough spheres to the fibres of the deformation $F_{t}$ above an appropriate small interval at 0. This derives from a principle established by Milnor [Mi] and which occurs, mostly under the name of $\rho$-regularity, in several studies of fibrations of real analytic map germs on smooth or singular base spaces, such as [ACT1], [ACT2], [AT1], [AT2], [ART], [JoT1] etc, and see [Ti2] for older references.
@@
Let $F_{t}(x)=F(x,t)$ be a C1-family of analytic function germs $F_{t}:({\mathbb{K}}^{n},0)\to({\mathbb{K}},0)$. We consider the following condition:
(3) $\quad&\mbox{For any }x_{0}\in{\mathrm{Sing\hskip 2.0pt}}F_{0}\setminus\{0\},\mbox{ there is }c_{x_{0}}>0\mbox{ such that }\\ \quad&\biggl{|}\frac{\partial F}{\partial t}\biggr{|}\leq c_{x_{0}}\biggl{\|}\frac{\partial F}{\partial x_{1}},\ldots,\frac{\partial F}{\partial x_{n}}\biggr{\|}\ \mbox{ when }(x,t)\to(x_{0},0).\end{aligned}\right.$
@@
We prove here:
@@
Then the deformation $F$ of $F_{0}$ is topologically trivial at the boundary in the sense of Definition 1.1. Moreover, $\widetilde{F}$ has a Milnor-Hamm tube fibration (6), cf Definition 3.1.
The proof follows from Theorem 3.7 after showing by Theorem 4.3 that a nice deformation is tame (Definition 3.5). The tameness follows from the proof of the existence of the Milnor-Hamm tube fibration which is based on checking the Thom regularity condition, a well-known method employed e.g. in [Hi], [LS], [Sa1, Sa2], [PT], [AT2], [ACT1] etc. Here we use a less demanding condition called $\partial$-Thom regularity (see §4.1 for details), and that the $\partial$-Thom regularity implies the $\rho$-regularity (Proposition 4.2).
Our second condition is algebraic. We show by Theorem 5.6 that if the Jacobian ideal inclusion $(\partial F_{t})\subset(\partial F_{0})$ holds for any $t$ close enough to 0, then condition (3) holds, hence the deformation $F$ is tame, and thus it is “topologically trivial at the boundary” by our general Theorem 4.3. To a certain extent, this comes closer to Manolescu’s Question 1. Moreover, in §5 we slightly extend the setting by replacing the ideal $(\partial F_{0})$ in the above inclusion by its integral closure $\overline{(\partial F_{0})}$.
@@
Acknowledgements. We thank Ciprian Manolescu for bringing to our attention Question 1. We also thank him and Laurenţiu Păunescu for enriching observations which impelled our study.
@@ 2. Image and discriminant of a deformation
$\widetilde{F}:=(F(x,t),t):({\mathbb{K}}^{n}\times{\mathbb{K}},0)\to({\mathbb{K}}\times{\mathbb{K}},0)$
be the analytic map defined by this deformation. The singular set ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$ is the zero locus $Z\bigl{(}\frac{\partial F}{\partial x_{1}}(x,t),\ldots,\frac{\partial F}{\partial x_{n}}(x,t)\bigr{)}$. As a set germ at the origin $(0,0)$, it contains the union of set germs $\bigcup_{t\in{\mathbb{K}}}{\mathrm{Sing\hskip 2.0pt}}F_{t}$, but may contain other irreducible components. Note that ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}\cap\{t=0\}={\mathrm{Sing\hskip 2.0pt}}F_{0}$, and that we have the inclusion ${\mathrm{Sing\hskip 2.0pt}}F\subset{\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$, which may be a strict inclusion even when restricted to the slice $\{t=0\}$, like in the example: $F(x,y,t)=x^{2}+ty$.
@@ 2.1. The image problem
In case ${\mathbb{K}}={\mathbb{C}}$ the map $\widetilde{F}$ is locally open.
@@
    If $F_{0}$ is locally open then $\widetilde{F}$ is a locally open map. In particular this is the case if ${\mathrm{Sing\hskip 2.0pt}}F_{0}\not=F_{0}^{-1}(0)$.
@@ 3.1. Local tube fibration
We say that $\widetilde{F}$ has a local tube fibration, also called Milnor-Hamm fibration, if for any small enough $r>0$ there exist $0<\eta\ll\delta\ll r$ such that the restriction:
@@
is a locally trivial fibration which is independent of the chosen constants, up to isotopy.
@@
If the local tube fibration (6) exists, then the deformation $F(x,t)$ is topologically trivial at the boundary (in the sense of Definition 1.1).
@@
If the analytic deformation $F(x,t)$ of $F(x,0)$ is tame, then $\widetilde{F}$ has a Milnor-Hamm tube fibration (6), and the deformation $F(x,t)$ is topologically trivial at the boundary (in the sense of Definition 1.1).
@@
We have seen that in case of a deformation $F$, the condition “nice” holds in the holomorphic setting, thus the above result applies and yields our claim.
@@ 4.1. The partial Thom stratification
Let us recall the $\partial$-Thom regularity, after [Ti1, Def. 2.1], [Ti2, A 1.1], [DRT, §6], and [ART, §4]. This is a weaker condition than the usual Thom regularity, but sufficient for the existence of the Milnor tube fibration in the general case of real map germs, see Proposition 4.2 below. There are however examples where the map germ $G$ has Milnor tube fibration without being $\partial$-Thom regular, see [PT] and [ART].
@@ 4.2. Theorem and its proof
(10) $\lim_{(x_{t},t)\to(y,0)}T_{(x_{t},t)}F^{-1}(s_{t})\supset T_{(y,0)}V.$
@@
where $V^{\prime}$ is the stratum of ${\mathcal{S}}^{\prime}$ which contains $(y,0)$, in particular we have $T_{(y,0)}V\supset T_{(y,0)}V^{\prime}$.
We will deduce it from the inclusion (10) via the condition (3). Let us suppose by reductio ad absurdum that we have:
@@
The existence of the partial Thom stratification ${\mathcal{S}}^{\prime}$ for the map $\widetilde{F}$ implies that there is $R>0$ such that, for any positive $r\leq R$, the sphere $S_{r}\subset{\mathbb{K}}^{n}$ is transversal to all positive dimensional strata of ${\mathcal{S}}^{\prime}$. As a direct consequence of the definition (10), it follows that the sphere $S_{r}$ is transversal to the smooth nearby fibres of $\widetilde{F}$ as ine (6), and therefore $F$ is tame (and in particular, by Theorem 3.7, the tube fibration (6) exits). This ends the proof of Theorem 4.3. ∎
@@ 5. The Jacobian criterium
The integral closure of an ideal $J\subset{\mathcal{A}}_{n}$, denoted by $\overline{J}$, is the set of all $f\in{\mathcal{A}}_{n}$ such that for any real analytic arc $\mu:({\mathbb{R}},0)\to({\mathbb{R}}^{n},0)$ one has $\mu^{*}f\in(\mu^{*}J){\mathcal{A}}_{1}$, equivalently:
@@
In the complex analytic setting, it was proved by Teissier [Te1] that the usual algebraic definition of the integral closure is equivalent to the above one. In the real analytic case, the above definition was introduced by Gaffney [Ga].
@@
Then the deformation $F$ of $F_{0}$ is topologically trivial at the boundary in the sense of Definition 1.1, and $\widetilde{F}$ has a Milnor-Hamm tube fibration (6), cf Definition 3.1.
@@
Any real analytic function germ extends uniquely to a holomorphic function germ (in terms of representatives, the same function can be viewed as a holomorphic function on a complex domain). Let us denote by $I_{{\mathbb{C}}}\subset{\mathcal{O}}_{n}$ the ideal of holomorphic function germs generated by the holomorphic extensions of the generators of $I$. We follow the same pattern as above, namely we apply Cartan’s Theorem to (18) but for the complex ideal $I_{{\mathbb{C}}}$ instead of the real ideal $I$. Like done just above, we then deduce that $\partial_{i}f_{1}\in I_{{\mathbb{C}}}$. This means that $\partial_{i}f_{1}$ is equal to a linear combination with complex analytic coefficients of the real generators of $I_{{\mathbb{C}}}$. But since (the restriction to ${\mathbb{R}}^{n}$ of) $\partial_{i}f_{1}$ is a real function, it follows that the complex part of that linear combination is zero. Thus by taking the real part of each coefficient yields a real linear combination equal to $\partial_{i}f_{1}$. This shows that we actually have $\partial_{i}f_{1}\in I$, and the proof of point (a) is done.
To show (b) we proceed in the same manner, we interpret the real analytic functions as complex analytic and apply Cartan’s theorem as explained above to get $\partial_{i}\bigl{(}\frac{\partial F}{\partial t}\bigr{)}\in I_{{\mathbb{C}}}$. Now, by the same arguments for the linear combination, we get that actually $\partial_{i}\bigl{(}\frac{\partial F}{\partial t}\bigr{)}\in I$. ∎
@@
Theorem 5.8 is not true over ${\mathbb{R}}$. A counter-example can be found for instance in [ABF, Ch. 4 §6]. Namely, one constructs a sequence of real functions $\{f_{k}\}$ which are analytic on the real line ${\mathbb{R}}$, that converge uniformly on compacts to some analytic function $f$, and such that the germ at the origin of each $f_{k}$ is in some ideal $I$, but that $f$ is not in $I$. The problem is that the radius of convergency of the Taylor series of $f_{k}$ at the origin goes to zero as $k\to\infty$. In contrast, note that the setting of our above proof insures the existence of a fixed small neighbourhood of the origin in which the Taylor series of $f_{j}$ converges, for all $j\geq 0$.
@@
Let us show that condition (16) is satisfied. We apply Proposition 5.3 to the ideal $(\partial F_{0})$ and its generators $\partial_{1}F_{0},\dots,\partial_{2}F_{0}$, and for the choice of a neighbourhood $U$ of the origin such that the conclusion of Proposition 5.3 holds.
@@
By Proposition 5.5 it then follows that condition (3) is satisfied, and thus by Theorem 4.3 and Theorem 3.7, it follows that $F$ is tame, that $\widetilde{F}$ has a Milnor-Hamm tube fibration (6), and that $F$ is topologically trivial at the boundary (in the sense of Definition 1.1).

### v3

@@ Abstract.
We give analytic and algebraic conditions under which a deformation of real analytic functions with non-isolated singular locus is a deformation with fibre constancy.
@@
deformations, fibrations, fibre constancy
@@
The authors acknowledge the partial support of the Labex CEMPI (ANR-11-LABX-0007-01). C. Joiţa and M. Tibăr acknowledges support from the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan, and support by the grant CNRS-INSMI-IEA-329.
@@ 1. Introduction
In the complex setting, for holomorphic function germs with isolated singularities (i.e. the very special case of the 0-dimensional singular locus), the constancy of the Jacobian ideal $(\partial F_{t})$ obviously implies that the Milnor number $\mu(F_{t})=\dim_{{\mathbb{C}}}\frac{{\mathcal{O}}_{n}}{(\partial F_{t})}$ is constant, and thus the homology of the Milnor fibre is constant too since it is concentrated in the group $\tilde{H}_{n-1}(F_{t},{\mathbb{Z}})$, which is free of rank precisely $\mu(F_{t})$. Moreover, by the Lê-Ramanujam [LR], Timourian [Tim], and King [Ki] results, the constancy of the Milnor number implies that the deformation $F_{t}$ is topologically trivial with the exception of the surface case (i.e. $n=3$) which remains open.
In the case of complex non-isolated singularities, this question seems to be open in general. There is nevertheless a rich literature on equisingularity problems for families of function germs, dealing with various algebraic-geometric sufficient conditions, starting with the Whitney equisingularity studied by Teissier [Te1, Te2], extended by Gaffney [Ga], Houston [Hou] and many others.
In the real setting there are results by King [Ki] on the topological triviality of one-parameter families of function germs with isolated singularity. While the non-isolated case remains widely open in general, Parusinski’s paper [Pa2] goes beyond “isolated singularity” in the setting of families of the form $F(x,t)=f(x)+tg(x)$ subject to the condition:
@@
In this paper we are not interested in the problem of the topological triviality of families of function germs. Instead, we will answer Question 1 in a different way, by focussing on the constancy of the fibration (2) in a fixed ball, as explained in the following.
@@
Let $F:({\mathbb{K}}^{n}\times{\mathbb{K}},0)\to({\mathbb{K}}^{m},0)$, $n\geq m\geq 1$, ${\mathbb{K}}={\mathbb{R}}$ or ${\mathbb{C}}$, be a ${\mathbb{K}}$-analytic map germ, regarded as a 1-parameter deformation $F_{t}(x)=F(x,t)$ of the function germ $F_{0}=f:({\mathbb{K}}^{n},0)\to({\mathbb{K}},0)$ having a singular locus of dimension $\geq 1$.
We say that the deformation $F$ of the singular function germ $f$ is a deformation with fibre constancy if for any small enough radius $r>0$, there exist $0<\delta:=\delta(r)\ll r$, and $0<\eta:=\eta(r,\delta)\ll\delta$ such that the restriction:
@@
is a locally trivial fibration, which is independent, modulo isotopy, of $r$ and $\delta$, and of the parameter $t$ with $|t|\leq\eta\ll\delta$.
@@
In the real setting, $D_{\delta}$ denotes an interval centred at the origin and its boundary $\partial D_{\delta}=\{a_{-},a_{+}\}$ consists of two points. In this setting, the independency asked by Definition 1.1 means the invariance of the two fibres [^1] Of which one may be empty, but not both empty since we assume that the function germ $F_{0}$ is non-constant., $B\cap F_{t}^{-1}(a_{-})$ and $B\cap F_{t}^{-1}(a_{+})$. One may then ask:
@@
Question 2. What natural (and minimal) condition implies that the deformation $F$ is a deformation with fibre constancy?
@@
The problem posed by Question 2 is not a classical equisingularity problem and makes really sense in the case of $f$ with non-isolated singularities. If the function germ $F_{0}$ has an isolated singularity, then for $t\not=0$ this may either split into several isolated singularities of the restriction map $(F_{t})_{|}$ from (2) above, or not split at all. Even if the isolated singularity of $F_{0}$ splits, the tube fibrations (2) still exists precisely because the splitting is concentrated at the origin only (i.e. because of the stated choice of a much smaller range of the parameter $t$, namely $|t|\ll\delta$), see e.g. [JiT] for the study of a slightly more general setting of isolated singularities. In case it does not split, then one has stronger results: the above cited ones by Lê-Ramanujam [LR] and Timourian [Tim] over ${\mathbb{C}}$, and by King [Ki] over ${\mathbb{R}}$, yield the topological triviality of the family of function germs $F_{t}$, or what one calls topological equisingularity.
@@
In the setting of non-isolated holomorphic functions, the concept of “deformation with fibre constancy” (Definition 1.1) appeared under the name of “admissible deformations” in the studies of the topology of certain classes of function germs with non-isolated singular locus, starting with the seminal paper by Siersma [Si1] on “line singularities”, and the follow-up more general studies of the Milnor fibre of function germs with 1-dimensional and 2-dimension singular loci, see e.g. [Si2], [Pe], [dJ], [Za], [MS], [Ne], [Fe], [FM], [ST2]. The very recent paper [Hof] gives algebro-geometric criteria which are sufficient to establish admissibility for complex-analytic deformations.
@@
In both settings, complex and real, our general answer to Question 2 is (Theorem 3.7): “tame deformations” are deformations with fibre constancy. This “tameness” condition, introduced in Definition 3.5, is based on the transversality of all small enough spheres to the fibres of the deformation $(F_{t})_{|}$ defined at (2) above an appropriate small disk $D_{\delta}$. This derives from a principle established by Milnor [Mi] and which occurs, mostly under the name of $\rho$-regularity, in several studies of fibrations of real analytic map germs on smooth or singular base spaces, such as [ACT1], [ACT2], [AT1], [AT2], [ART], [JoT1], [CJT], and see [Ti2] for older references.
@@
Let $F_{t}(x)=F(x,t)$ be an analytic deformation of $f$ (cf Definition 1.1). We consider the following condition:
(3) $\quad&\mbox{For any }x_{0}\in{\mathrm{Sing\hskip 2.0pt}}F_{0}\setminus\{0\},\mbox{ there is }c_{x_{0}}>0\mbox{ such that }\\ \quad&\biggl{|}\frac{\partial F}{\partial t}\biggr{|}\leq c_{x_{0}}\biggl{\|}\frac{\partial F}{\partial x_{1}},\ldots,\frac{\partial F}{\partial x_{n}}\biggr{\|}\ \mbox{ when }(x,t)\to(x_{0},0),\mbox{ for }(x,t)\not\in{\mathrm{Sing\hskip 2.0pt}}\widetilde{F}.\end{aligned}\right.$
where $\widetilde{F}(x,t):=\bigl{(}F(x,t),t\bigr{)}$, where $\widetilde{F}(x,t)$ is the map germ defined at (5).
@@
Here we go beyond the isolated singularity setting, and prove the following:
@@
Then the deformation $F$ of $F_{0}$ is a deformation with fibre constancy in the sense of Definition 1.1. Moreover, $\widetilde{F}$ has a Milnor-Hamm tube fibration (7), cf Definition 3.1.
The proof follows from Theorem 3.7 after showing in Theorem 4.3 that a nice deformation is tame (Definition 3.5). The tameness follows from the proof of the existence of the Milnor-Hamm tube fibration which is based on checking the Thom regularity condition, a well-known method employed e.g. in [Hi], [LS], [Sa1, Sa2], [PT], [AT2], [ACT1] etc. Here we use a less demanding condition called $\partial$-Thom regularity (see §4.1 for details), and that the $\partial$-Thom regularity implies the $\rho$-regularity (Proposition 4.2).
Our second condition is algebraic. We show by Theorem 5.6 that if the Jacobian ideal inclusion $(\partial F_{t})\subset(\partial F_{0})$ holds for any $t$ close enough to 0, then condition (3) holds, hence the deformation $F$ is tame, and thus it is “a deformation with fibre constancy” by our general Theorem 4.3. To a certain extent, this comes closer to Manolescu’s Question 1. Moreover, in §5 we slightly extend the setting by replacing the ideal $(\partial F_{0})$ in the above inclusion by its integral closure $\overline{(\partial F_{0})}$.
@@
Acknowledgements. We thank Ciprian Manolescu for bringing to our attention Question 1. We also thank him and Laurenţiu Păunescu for enriching observations which impelled our study. Special thanks go to the referee for very interesting remarks which helped us improve the paper.
C. Joiţa and M. Tibăr acknowledge support of the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan, and support by the grant CNRS-INSMI-IEA-329.
@@ 2. Image and discriminant of a deformation
(5) $\widetilde{F}:=(F(x,t),t):({\mathbb{K}}^{n}\times{\mathbb{K}},0)\to({\mathbb{K}}\times{\mathbb{K}},0)$
be the analytic map germ defined by the deformation $F$. The singular set ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$ is the zero locus $Z\bigl{(}\frac{\partial F}{\partial x_{1}}(x,t),\ldots,\frac{\partial F}{\partial x_{n}}(x,t)\bigr{)}$. As a set germ at the origin $(0,0)$, it contains the union of set germs $\bigcup_{t\in{\mathbb{K}}}{\mathrm{Sing\hskip 2.0pt}}F_{t}$, but may contain other irreducible components, see e.g. Example 6.1. Note that ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}\cap\{t=0\}={\mathrm{Sing\hskip 2.0pt}}F_{0}$, and that we have the inclusion ${\mathrm{Sing\hskip 2.0pt}}F\subset{\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$, which may be a strict inclusion even when restricted to the slice $\{t=0\}$, like in the example: $F(x,y,t)=x^{2}+ty$.
@@ 2.1. The image problem
In case ${\mathbb{K}}={\mathbb{C}}$, the map germ $\widetilde{F}$ is locally open.
@@
    If $F_{0}$ is locally open then $\widetilde{F}$ is a locally open map germ. In particular this is the case if ${\mathrm{Sing\hskip 2.0pt}}F_{0}\not=F_{0}^{-1}(0)$.
@@ 3.1. Local tube fibration
We say that $\widetilde{F}$ has a local tube fibration, also called Milnor-Hamm fibration, if for any small enough $r>0$ there exists a radius $\delta\ll r$, and a radius $\eta\ll\delta$ such that the restriction:
@@
is a locally trivial fibration which is independent of the chosen constants up to isotopy.
@@
If the local tube fibration (7) exists, then the deformation $F(x,t)$ is a deformation with fibre constancy (in the sense of Definition 1.1).
@@
If the analytic deformation $F(x,t)$ of $F(x,0)$ is tame, then $\widetilde{F}$ has a Milnor-Hamm tube fibration (7), and the deformation $F(x,t)$ is a deformation with fibre constancy (in the sense of Definition 1.1).
@@
We have seen in §2.1 that in case of a deformation $F$, the condition “nice” (recalled after Definition 3.1) holds in the holomorphic setting, thus the above result applies and yields our claim.
@@ 4.1. The partial Thom stratification
Let us recall the $\partial$-Thom regularity, after [Ti1, Def. 2.1], [Ti2, A 1.1], [DRT, §6], and [ART, §4], see also [CJT]. This is a weaker condition than the usual Thom regularity, but sufficient for the existence of the Milnor tube fibration (cf Definition 3.1) in the general case of real map germs, see Proposition 4.2 below. There are however examples where the map germ $G$ has Milnor tube fibration without being $\partial$-Thom regular, see [PT] and [ART].
@@ 4.2. Theorem and its proof
(11) $\lim_{(x_{t},t)\to(y,0)}T_{(x_{t},t)}F^{-1}(s_{t})\supset T_{(y,0)}V,$
where we may assume without loss of generality that the limit exists in the appropriate Grassmannian.
@@
where by $V^{\prime}$ we denote the stratum of ${\mathcal{S}}^{\prime}$ which contains $(y,0)$, in particular we have $T_{(y,0)}V\supset T_{(y,0)}V^{\prime}$.
We will deduce (12) from the inclusion (11) via the condition (3). Let us suppose by reductio ad absurdum that we have:
@@
The existence of the partial Thom stratification ${\mathcal{S}}^{\prime}$ for the map $\widetilde{F}$ implies that there is $R>0$ such that, for any positive $r\leq R$, the sphere $S_{r}\subset{\mathbb{K}}^{n}$ is transversal to all positive dimensional strata of ${\mathcal{S}}^{\prime}$. As a direct consequence of the definition (11), it follows that the sphere $S_{r}$ is transversal to the smooth nearby fibres of $\widetilde{F}$ as in (7), and therefore $F$ is tame (and in particular, by Theorem 3.7, the tube fibration (7) exits). This ends the proof of Theorem 4.3. ∎
@@ 5. The Jacobian criterium
The integral closure of an ideal $J\subset{\mathcal{A}}_{n}$, denoted by $\overline{J}$, is the set of all $f\in{\mathcal{A}}_{n}$ such that for any analytic arc $\mu:({\mathbb{K}},0)\to({\mathbb{K}}^{n},0)$ one has $\mu^{*}f\in(\mu^{*}J){\mathcal{A}}_{1}$, equivalently:
@@
The above definition was introduced by Gaffney [Ga] in the real analytic setting. In the complex analytic setting, it was proved by Teissier [Te1] that the usual algebraic definition of the integral closure is equivalent to the above one.
@@
Then the deformation $F$ of $F_{0}$ is a deformation with fibre constancy in the sense of Definition 1.1, and $\widetilde{F}$ has a Milnor-Hamm tube fibration (7), cf Definition 3.1.
@@
We choose a polydisk $\{(x,t)\in{\mathbb{C}}^{n}\times{\mathbb{C}}\mid|x_{i}|<\alpha_{i},|t|<\delta,i=1,\dots,n\}$ on which $F$ is holomorphic, i.e. for some well-chosen constants $\alpha_{i}>0$. Hence each $f_{j}$ is holomorphic on $V:=\{x\in{\mathbb{C}}^{n}\mid|x_{i}|<\alpha_{i},i=1,\dots,n\}$.
@@
We choose $\delta>0$, $\alpha_{i}>0$, $i\in\{1,\dots,n\}$, such that the Taylor series of $F$ at $(0,0)\in{\mathbb{R}}^{n}\times{\mathbb{R}}$ converges on $\{|x_{i}|<\alpha_{i}\mid i=1,\dots,n\}\times\{|t|<\delta\}$. In particular the Taylor series of each $f_{j}$ converges on $\{|x_{i}|<\alpha_{i}\mid i=1,\dots,n\}\subset{\mathbb{R}}^{n}$.
Let $F^{c}$ be the complexification of $F$ (i.e. if $\sum_{J\subset{\mathbb{N}}^{n},j\in{\mathbb{N}}}a_{j,J}t^{j}x^{J}$ is the Taylor series of $F$ then $F^{c}(z,\lambda)=\sum_{J\subset{\mathbb{N}}^{n},j\in{\mathbb{N}}}a_{j,J}\lambda^{j}z^{J}$ for $z_{j},\lambda\in{\mathbb{C}}$) and let $f^{c}_{j}$ be the complexification of $f_{j}$.
We consider now the polydisk $P\subset{\mathbb{C}}^{n}\times{\mathbb{C}}$, $P=\{(z,\lambda)\mid|z_{i}|<\alpha_{i},|\lambda|<\delta\}$. Then $F^{c}$ is holomorphic on $P$. Let $I_{{\mathbb{C}}}$ be the ideal in ${\mathcal{O}}_{n}$ generated by $\partial_{z}F^{c}_{0}$.
If $t\in{\mathbb{R}}$, $|t|<\delta$, is such that $\langle\partial_{x}F_{t}\rangle\subset I$, then it follows that $\langle\partial_{z}F^{c}_{t}\rangle\subset I_{{\mathbb{C}}}$. We notice now that in the proof of the ${\mathbb{C}}$-analytic case, we did not actually need the inclusion of the Jacobian ideals for all complex values of the parameter, namely some sequence convergent to 00 suffices. We obtain that $\partial f_{j}^{c}\subset I_{{\mathbb{C}}}$. This means that $\partial_{i}f^{c}_{j}$ is a linear combination with complex analytic coefficients of the generators of $I_{{\mathbb{C}}}$. But since the restriction to ${\mathbb{R}}^{n}$ of $\frac{\partial f^{c}_{j}}{\partial z_{i}}$ is $\frac{\partial f_{j}}{\partial x_{i}}$, hence a real function, it follows that, by taking the real part of each coefficient, we obtain a real linear combination equal to $\partial_{i}f_{j}$. This shows that we actually have $\partial_{i}f_{j}\in I$, and the proof of point (a) is done.
To show (b) we proceed in the same manner: we interpret the real analytic functions as complex analytic and we apply Cartan’s theorem, as explained above, to get $\partial_{i}\bigl{(}\frac{\partial F}{\partial t}\bigr{)}\in I_{{\mathbb{C}}}$. Now, by the same arguments for the linear combination, we get that actually $\partial_{i}\bigl{(}\frac{\partial F}{\partial t}\bigr{)}\in I$. ∎
@@
Theorem 5.8 is not true over ${\mathbb{R}}$. A counter-example can be found for instance in [ABF, Ch. 4 §6]. Namely, one constructs a sequence of real functions $\{f_{k}\}$ which are analytic on the real line ${\mathbb{R}}$, that converge uniformly on compacts to some analytic function $f$, and such that the germ at the origin of each $f_{k}$ is in some ideal $I$, but that $f$ is not in $I$. The problem is that the radius of convergency of the Taylor series of $f_{k}$ at the origin goes to zero as $k\to\infty$. In contrast, note that the setting of our above proof insures a stronger convergence, as defined e.g. in [ABF, Definition 6.2].
@@
Let us show that condition (17) is satisfied. We apply Proposition 5.3 to the ideal $(\partial F_{0})$ and its generators $\partial_{1}F_{0},\dots,\partial_{n}F_{0}$, and for the choice of a neighbourhood $U$ of the origin such that the conclusion of Proposition 5.3 holds.
@@
By Proposition 5.5 it then follows that condition (3) is satisfied, and thus by Theorem 4.3 and Theorem 3.7, it follows that $F$ is tame, that $\widetilde{F}$ has a Milnor-Hamm tube fibration (7), and that $F$ is a deformation with fibre constancy (in the sense of Definition 1.1).